OpenStudy (anonymous):

2/3+2/3=? I will give medal.

OpenStudy (anonymous):

@KamiBug and @Secret-Ninja

OpenStudy (secret-ninja):

\[\frac{ 2 }{ 3 } + \frac{ 2 }{ 3 }\] Like that?

OpenStudy (anonymous):

switch the sign and flip 2/3 to 3/2. then add

OpenStudy (secret-ninja):

Well, since the denominators are already the same, they stay at 3 and you just add the 2 +2 which is 4 so you get \[\frac{ 4 }{ 3 }\]

OpenStudy (secret-ninja):

@DraymondBrown Is incorrect. That is not for addition.

OpenStudy (secret-ninja):

When you take the 4/3 and turn it into a mixed fraction you get \[1 \frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

ohhhhhhhhhhhhh haha thx @Secret-Ninja

OpenStudy (secret-ninja):

No problem ;]